A SAMPLING METHOD FOR AN INVERSE BOUNDARY VALUE PROBLEM FOR HARMONIC VECTOR FIELDS

RAINER KRESS

Abstract. Two variants of sampling methods have been recently suggested and developed by Colton and Kirsch for the approximate solution of inverse obstacle scattering problems for acoustic waves, i.e., for inverse boundary value problems for the Helmholtz equation. We present an analogue of the second variant of these methods for the solution of an inverse boundary value problem from electrostatics, i.e., for an inverse Dirichlet problem for harmonic vector fields.